Determine the resultant displacement of the superposition of the following harmonic waves

y1 = 7sin(wt + π/3), y2 = 12cos(wt + π/4) and y3 = 20sin(wt + π/5)


1
Expert's answer
2021-02-03T02:47:24-0500

Answer

Resultant of y1 and y2

"y=y_1+y_3=26.5\sin(wt-18^°)"

Now resultant of y and y2

"R=y+y_2"

="23.68\sin(wt+5^°)"